OpenStudy (anonymous):

Solve the equation for y. Then find the value of y for each value of x. Please help me, and thanks! 5x = -4y + 4; x = 1, 2, 3

OpenStudy (anonymous):

can you solve for y in this equation: 5x = -4y + 4 ????

OpenStudy (anonymous):

idk. :/

OpenStudy (anonymous):

5x = -4y + 4 5x -4 = -4y + 4 -4 subtract 4 from both sides 5x - 4 = -4y simplify \(\large \frac{5x -4}{-4}= \frac{-4y}{-4} \) divide -4 to both sides... \(\large \frac{4-5x}{4}= y \) simplify now just plug those values for x....